# Error Bounds and Optimal Schedules for Masked Diffusions with Factorized Approximations

### v1

@@ 1 Set-up, background and objectives
Assume we are interested in generating samples from a probability distribution $\pi$ on a product space $\mathcal{X}^{N}$, where $\mathcal{X}$ can be a finite set of tokens or some other general state space. A standard approach to generate a sample $\mathbf{x}=(x_{1},\ldots,x_{N})$ is to proceed sequentially, sampling first $x_{1}$ from its marginal distribution $\pi(x_{1})$ and then $x_{i}$ from its conditional distribution $\pi(x_{i}|\mathbf{x}_{<i})$, for $i=2,\dots,N$, where $\mathbf{x}_{<i}=(x_{1},\dots,x_{i-1})$. One limitation of this approach, underlying popular auto-regressive generative models (ARMs), is the need to perform $N$ sequential steps, which limits the speed and computational efficiency of the resulting algorithms. This has motivated the exploration of alternative procedures to generate samples from $\pi$ given access to its conditional distributions (or approximations thereof), such as masked diffusion models (MDMs) [1, 6, 15, 14]. Despite being originally derived by analogy with diffusion models on continuous state spaces, MDMs can also be understood as a way to reduce the cost of standard ARMs by generating multiple tokens simultaneously, see e.g. [8, Sec. 2.1.2] and references therein for more discussion. In this work, we follow this more algorithmic perspective, directly defining a general class of ‘unmasking’ (or ‘sequential sampling’) algorithms and analysing their properties and the resulting computational-vs-accuracy trade-off.
@@
where the sum runs over $z$ subset of $\{1,\ldots,N\}$ and $i\notin z$, with weights $\omega(i,z)$ which typically are chosen to be uniform; see e.g. [16, Sec. 3] and [15]. By construction, this loss is minimized by the exact conditionals of $\pi$. Models of $p^{\theta}$ used in practice receive $\mathbf{x}_{\mathbf{z}_{<k}}$ as input and produce all univariate conditional distributions $\{p^{\theta}(x_{i};\mathbf{x}_{\mathbf{z}_{<k}})\}_{i\notin\mathbf{z}_{<k},x_{i}\in\mathcal{X}}$ as output. Once the model $p^{\theta}$ has been evaluated, the cost of sampling from the $|z_{k}|$ univariate conditionals in (2) is comparatively small. As a result, the dominant cost in Algorithm 1 is given by the $K$ evaluations of the model $p^{\theta}$, which is why we measure the computational cost of Algorithm 1 with $K$. See e.g. the discussion about sampling efficiency of MDMs in Ben-Hamu et al. [4], where $K$ is referred to as number of function evaluations.
@@ 2 Error decomposition
We will use the monotonicity of the Kullback-Leibler divergence: the $\mathrm{KL}$ decreases if we marginalize, and it also decreases (in expected value) if we condition. Both properties come from the chain rule for $\mathrm{KL}$ [7, Theorem 2.5.3].
@@
$\mathrm{TC}_{\pi}(z_{k}|\mathbf{z}_{<k})=\mathbb{E}_{\pi(\mathbf{x})}\left[\log\frac{\pi(\mathbf{x}_{z_{k}}|\mathbf{x}_{\mathbf{z}_{<k}})}{\prod_{i\in z_{k}}\pi(x_{i}|\mathbf{x}_{\mathbf{z}_{<k}})}\right]\,.$
It measures how correlated the components of $x_{i}$, $i\in z_{k}$ are, given $x_{j}$, $j\in\mathbf{z}_{<k}$, see e.g. [2, Sec. 4] and references therein. By construction, $\mathrm{TC}(z_{k}|\mathbf{z}_{<k})$ is non-negative and it equals 0 if and only if the $x_{i}$, $i\in z_{k}$ are independent given $\mathbf{x}_{\mathbf{z}_{<k}}$, in particular this is always the case if $z_{k}$ contains exactly one element.
@@
$E_{\text{learn}}=\mathbb{E}_{\pi(\mathbf{x})\nu^{\theta}(\mathbf{z};\mathbf{x})}\left[\sum_{k\geq 1}\sum_{i\in z_{k}}\log\frac{\pi(x_{i}|\mathbf{x}_{\mathbf{z}_{<k}})}{p^{\theta}(x_{i};\mathbf{x}_{\mathbf{z}_{<k}})}\right]\,,\qquad E_{\text{fact}}=\mathbb{E}_{\pi(\mathbf{x})\nu^{\theta}(\mathbf{z};\mathbf{x})}\left[\sum_{k\geq 1}\mathrm{TC}_{\pi}(z_{k}|\mathbf{z}_{<k})\right]\,.$
@@
Re-arranging the last two equations gives $\mathrm{KL}({\pi(\mathbf{x})\nu^{\theta}(\mathbf{z};\mathbf{x})}\|{p^{\theta}(\mathbf{x};\mathbf{z})\nu^{\theta}(\mathbf{z};\mathbf{x})})=E_{\text{learn}}+E_{\text{fact}}$. ∎
@@ 3 Worst-case bounds on the factorization error
Moreover, denoting $s_{k}=|z_{k}|$ and $s_{\max}=\max(s_{1},\ldots,s_{K})$, we have
$E_{\text{fact}}\leq\left(N-\mathbb{E}_{\pi(\mathbf{x})\nu^{\theta}(\mathbf{z};\mathbf{x})}\left[\frac{N}{s_{\max}}\right]\right)D(\pi)\,.$ (5)
We use the following decomposition for the total correlation: for any ordering $i_{1},\ldots i_{s_{k}}$ of $z_{k}$, as $\pi(\mathbf{x}_{z_{k}}|\mathbf{x}_{\mathbf{z}_{<k}})=\prod_{\ell=1}^{s_{k}}\pi(x_{i_{\ell}}|\mathbf{x}_{\mathbf{z}_{<k}\cup\{i_{1},\ldots i_{\ell-1}\}})$ we have
$\mathrm{TC}_{\pi}(z_{k}|\mathbf{z}_{<k})=\sum_{\ell=2}^{s_{k}}\mathbb{E}_{\pi(\mathbf{x})}\left[\log\frac{\pi(x_{i_{\ell}}|\mathbf{x}_{\mathbf{z}_{<k}\cup\{i_{1},\ldots i_{\ell-1}\}})}{\pi(x_{i_{\ell}}|\mathbf{x}_{\mathbf{z}_{<k}})}\right]\leq\sum_{\ell=2}^{s_{k}}\mathbb{E}_{\pi(\mathbf{x})}\left[\log\frac{\pi(x_{i_{\ell}}|\mathbf{x}_{-i_{\ell}})}{\pi(x_{i_{\ell}})}\right],$
@@
$\mathrm{TC}_{\pi}(z_{k}|\mathbf{z}_{<k})\leq\left(1-\frac{1}{s_{k}}\right)\sum_{i\in z_{k}}\mathbb{E}_{\pi(\mathbf{x})}\left[\log\frac{\pi(x_{i}|\mathbf{x}_{-i})}{\pi(x_{i})}\right].$ (6)
Since $\mathbb{E}_{\pi(\mathbf{x})}[\log\pi(x_{i}|\mathbf{x}_{-i})/\pi(x_{i})]\leq-\mathbb{E}_{\pi(\mathbf{x})}[\log\pi(x_{i})]\leq\log|\mathcal{X}|$, we obtain $\mathrm{TC}_{\pi}(z_{k}|\mathbf{z}_{<k})\leq(s_{k}-1)\log|\mathcal{X}|$ and inequality (4) follows by summing over $k$ and using $\sum_{k}s_{k}=N$.
Finally, plugging (6) into the definition of $E_{\text{fact}}$, and using $\sum_{k}\sum_{i\in\mathbf{z}_{k}}\mathbb{E}_{\pi(\mathbf{x})}\left[\log\frac{\pi(x_{i}|\mathbf{x}_{-i})}{\pi(x_{i})}\right]=ND(\pi)$ for any partition $\mathbf{z}$, as well as $s_{k}\leq s_{\max}$, we obtain (5). ∎
@@
In this case a direct computation yields $\mathrm{TC}_{\pi}(z_{k}|\mathbf{z}_{<k})=(s_{k}-1)\log|\mathcal{X}|$, which gives equality in (4) by summing over $k$. Moreover in this case $\pi(x_{i}|\mathbf{x}_{-i})/\pi(x_{i})=1/\pi(x_{i})=|\mathcal{X}|$ and thus $D(\pi)=\log|\mathcal{X}|$. If in addition $N/s_{\max}=K$, it implies that there is also equality in (5). ∎
If $s_{k}$ is roughly constant over $k$, then $s_{\max}\approx N/K$, so that the upper bound in (5) is approximately $(N-K)D(\pi)$. In particular if $K,N\to+\infty$ and $N/K\to\bar{s}>1$, the bound in (4) grows linearly with $N$, resulting in very weak guarantees on the overall sampling error. However, such worst-case bounds can be very pessimistic: below we show that if $\mathbf{z}$ is randomized, then $E_{\text{fact}}$ is provably of much smaller size.
@@ 4 The random-order case
Consider now the situation where $z_{k}$ is sampled by first generating its size $s_{k}=|z_{k}|$ and then sampling its entries uniformly without replacement from $\{1,\dots,N\}\backslash\mathbf{z}_{<k}$. This is the case in common implementations of MDMs [15]. Specifically, to generate the ordered partition $\mathbf{z}$, the algorithm first sample the sizes $(s_{1},s_{2},\ldots,s_{K})=\mathbf{s}$ with $\sum_{k=1}^{K}s_{K}=N$. Then once these sizes are sampled, recursively
@@ 4.1 Rewriting the factorization error with the information profile
On the other hand, by (7), if $\mathbf{z}\sim\nu^{\theta}$ and $i\in z_{k}$ uniformly at random then $(i,\mathbf{z}_{<k})$ has distribution equal to $(\sigma_{a_{k-1}+1},\{\sigma_{1},\ldots\sigma_{a_{k-1}}\})$ with $\sigma\sim\mathrm{Unif}$. Thus as $|z_{k}|=s_{k}=a_{k}-a_{k-1}$
@@ 4.2 Resulting upper bounds
A bound similar to the one in Theorem 7 was recently derived in Li and Cai [9, Theorem 1] with a different and significantly less direct proof approach.
@@ 4.3 Explicit computations with geometric schedules and lower bounds
Interestingly, one can compute almost exactly the value of $E_{\text{fact}}$ for the case of random, geometrically distributed schedule sizes, with no assumption on $\pi$. Specifically, given $p\in(0,1)$ and $m\in\mathbb{N}$, let $\mathrm{Geom}(p;m)$ denote a Geometric distribution starting from $1$ and with a threshold at $m$, i.e. a random variable $X\sim\mathrm{Geom}(p;m)$ satisfies $\mathrm{Pr}(X=i)=(1-p)^{i-1}p$ for $i\in\{1,\dots,m-1\}$ and $\mathrm{Pr}(X=m)=(1-p)^{m-1}$. The proof of the following result relies on the memoryless property of the geometric distribution, it can be found in the appendix.
@@
$E_{\text{fact}}\leq\frac{1-p}{p}D(\pi),$ (12)
@@
$E_{\text{fact}}\geq\frac{1-p}{p}D(\pi)-\left(\frac{1-p}{p}\right)^{2}\left(\max_{i=1,\dots,N-1}\Delta f(i)\right)\,.$ (13)

### v2

@@ 1 Set-up, background and objectives
Assume we are interested in generating samples from a probability distribution $\pi$ on a product space $\mathcal{X}^{N}$, where $\mathcal{X}$ can be a finite set of tokens or some other general state space. A standard approach to generate a sample $\mathbf{x}=(x_{1},\ldots,x_{N})$ is to proceed sequentially, sampling first $x_{1}$ from its marginal distribution $\pi(x_{1})$ and then $x_{i}$ from its conditional distribution $\pi(x_{i}|\mathbf{x}_{<i})$, for $i=2,\dots,N$, where $\mathbf{x}_{<i}=(x_{1},\dots,x_{i-1})$. One limitation of this approach, underlying popular auto-regressive generative models (ARMs), is the need to perform $N$ sequential steps, which limits the speed and computational efficiency of the resulting algorithms. This has motivated the exploration of alternative procedures to generate samples from $\pi$ given access to its conditional distributions (or approximations thereof), such as masked diffusion models (MDMs) [1, 6, 16, 15]. Despite being originally derived by analogy with diffusion models on continuous state spaces, MDMs can also be understood as a way to reduce the cost of standard ARMs by generating multiple tokens simultaneously, see e.g. [9, Sec. 2.1.2] and references therein for more discussion. In this work, we follow this more algorithmic perspective, directly defining a general class of ‘unmasking’ (or ‘sequential sampling’) algorithms and analysing their properties and the resulting computational-vs-accuracy trade-off.
@@
where the sum runs over $z$ subset of $\{1,\ldots,N\}$ and $i\notin z$, with weights $\omega(i,z)$ which typically are chosen to be uniform; see e.g. [17, Sec. 3] and [16]. By construction, this loss is minimized by the exact conditionals of $\pi$. Models of $p^{\theta}$ used in practice receive $\mathbf{x}_{\mathbf{z}_{<k}}$ as input and produce all univariate conditional distributions $\{p^{\theta}(x_{i};\mathbf{x}_{\mathbf{z}_{<k}})\}_{i\notin\mathbf{z}_{<k},x_{i}\in\mathcal{X}}$ as output. Once the model $p^{\theta}$ has been evaluated, the cost of sampling from the $|z_{k}|$ univariate conditionals in (2) is comparatively small. As a result, the dominant cost in Algorithm 1 is given by the $K$ evaluations of the model $p^{\theta}$, which is why we measure the computational cost of Algorithm 1 with $K$. See e.g. the discussion about sampling efficiency of MDMs in Ben-Hamu et al. [4], where $K$ is referred to as number of function evaluations.
@@
Concurrent work
Similar results to ours were derived independently in the recent paper [7], which appeared online shortly after the first version of this work. In particular, [7] also relates the factorization error to the information profile of $\pi$ (see Result 1 and Theorem 1.4 therein) the same way our Lemma 6 does.
@@ 2 Error decomposition
We will use the monotonicity of the Kullback-Leibler divergence: the $\mathrm{KL}$ decreases if we marginalize, and it also decreases (in expected value) if we condition. Both properties come from the chain rule for $\mathrm{KL}$ [8, Theorem 2.5.3].
@@
$\mathrm{TC}_{\pi}(z_{k}|\mathbf{x}_{\mathbf{z}_{<k}})=\mathrm{KL}({\pi(\mathbf{x}_{z_{k}}|\mathbf{x}_{\mathbf{z}_{<k}})}\|{\otimes_{i\in z_{k}}\pi(x_{i}|\mathbf{x}_{\mathbf{z}_{<k}})})=\mathbb{E}_{\pi(\mathbf{x}_{z_{k}}|\mathbf{x}_{\mathbf{z}_{<k}})}\left[\log\frac{\pi(\mathbf{x}_{z_{k}}|\mathbf{x}_{\mathbf{z}_{<k}})}{\prod_{i\in z_{k}}\pi(x_{i}|\mathbf{x}_{\mathbf{z}_{<k}})}\right]\,$,
where $\otimes$ denotes the independent product of distributions. It measures how correlated the components of $x_{i}$, $i\in z_{k}$ are, given $x_{j}$, $j\in\mathbf{z}_{<k}$, see e.g. [2, Sec. 4] and references therein. By construction, $\mathrm{TC}_{\pi}(z_{k}|\mathbf{x}_{\mathbf{z}_{<k}})$ is non-negative and it equals 0 if and only if the $x_{i}$, $i\in z_{k}$ are independent given $\mathbf{x}_{\mathbf{z}_{<k}}$, in particular this is always the case if $z_{k}$ contains exactly one element.
@@
$E_{\text{learn}}=\mathbb{E}_{\pi(\mathbf{x})\nu^{\theta}(\mathbf{z};\mathbf{x})}\left[\sum_{k\geq 1}\sum_{i\in z_{k}}\log\frac{\pi(x_{i}|\mathbf{x}_{\mathbf{z}_{<k}})}{p^{\theta}(x_{i};\mathbf{x}_{\mathbf{z}_{<k}})}\right]\,,\qquad E_{\text{fact}}=\mathbb{E}_{\pi(\mathbf{x})\nu^{\theta}(\mathbf{z};\mathbf{x})}\left[\sum_{k\geq 1}\mathrm{TC}_{\pi}(z_{k}|\mathbf{x}_{\mathbf{z}_{<k}})\right]\,.$
@@
Taking the expectation with respect to $\pi(\mathbf{x})\nu^{\theta}(\mathbf{z};\mathbf{x})$ gives $\mathrm{KL}({\pi(\mathbf{x})\nu^{\theta}(\mathbf{z};\mathbf{x})}\|{p^{\theta}(\mathbf{x};\mathbf{z})\nu^{\theta}(\mathbf{z};\mathbf{x})})=E_{\text{fact}}+E_{\text{learn}}$ since the conditional distribution of $\mathbf{x}_{z_{k}}$ given $\mathbf{z}_{\leq k}$ and $\mathbf{x}_{\mathbf{z}_{<k}}$ under $\pi(\mathbf{x})\nu^{\theta}(\mathbf{z};\mathbf{x})$ coincides with $\pi(\mathbf{x}_{z_{k}}|\mathbf{x}_{\mathbf{z}_{<k}})$. ∎
@@ 3 Worst-case bounds on the factorization error
Moreover, if $\nu^{\theta}(\mathbf{z})=\nu^{\theta}(\mathbf{z};\mathbf{x})$ is independent of $\mathbf{x}$, denoting $s_{k}=|z_{k}|$ and $s_{\max}=\max(s_{1},\ldots,s_{K})$, we have
$E_{\text{fact}}\leq\left(N-\mathbb{E}_{\nu^{\theta}(\mathbf{z})}\left[\frac{N}{s_{\max}}\right]\right)D(\pi)\,.$ (5)
To prove (4), we use the following bound for the total correlation: for any ordering $i_{1},\ldots i_{s_{k}}$ of $z_{k}$, as $\pi(\mathbf{x}_{z_{k}}|\mathbf{x}_{\mathbf{z}_{<k}})=\prod_{\ell=1}^{s_{k}}\pi(x_{i_{\ell}}|\mathbf{x}_{\mathbf{z}_{<k}\cup\{i_{1},\ldots i_{\ell-1}\}})$ we have
$\mathrm{TC}_{\pi}(z_{k}|\mathbf{x}_{\mathbf{z}_{<k}})$ $=\sum_{\ell=2}^{s_{k}}\mathbb{E}_{\pi(\mathbf{x}_{z_{k}}|\mathbf{x}_{\mathbf{z}_{<k}})}\left[\log\frac{\pi(x_{i_{\ell}}|\mathbf{x}_{\mathbf{z}_{<k}\cup\{i_{1},\ldots i_{\ell-1}\}})}{\pi(x_{i_{\ell}}|\mathbf{x}_{\mathbf{z}_{<k}})}\right]$ (6)
$\leq-\sum_{\ell=2}^{s_{k}}\mathbb{E}_{\pi(\mathbf{x}_{z_{k}}|\mathbf{x}_{\mathbf{z}_{<k}})}\left[\log\pi(x_{i_{\ell}}|\mathbf{x}_{\mathbf{z}_{<k}})\right]\leq(s_{k}-1)\log|\mathcal{X}|\,.$
Summing over $k$ and using $\sum_{k}s_{k}=N$ gives $\sum_{k\geq 1}\mathrm{TC}_{\pi}(z_{k}|\mathbf{x}_{\mathbf{z}_{<k}})\leq(N-K)\log|\mathcal{X}|$, and then (4) follows from the definition of $E_{\text{fact}}$.
We now prove (5). If $\nu^{\theta}(\mathbf{z})=\nu^{\theta}(\mathbf{z};\mathbf{x})$ then $E_{\text{fact}}=\mathbb{E}_{\nu^{\theta}(\mathbf{z})}\left[\sum_{k\geq 1}\mathbb{E}_{\pi(\mathbf{x}_{z_{<k}})}[\mathrm{TC}_{\pi}(z_{k}|\mathbf{x}_{\mathbf{z}_{<k}})]\right]$. Taking expectations with respect to $\mathbf{x}_{z_{<k}}$ in (6), we obtain
$\mathbb{E}_{\pi(\mathbf{x}_{z_{<k}})}[\mathrm{TC}_{\pi}(z_{k}|\mathbf{x}_{\mathbf{z}_{<k}})]=\sum_{\ell=2}^{s_{k}}\mathbb{E}_{\pi(\mathbf{x})}\left[\log\frac{\pi(x_{i_{\ell}}|\mathbf{x}_{\mathbf{z}_{<k}\cup\{i_{1},\ldots i_{\ell-1}\}})}{\pi(x_{i_{\ell}}|\mathbf{x}_{\mathbf{z}_{<k}})}\right]\leq\sum_{\ell=2}^{s_{k}}\mathbb{E}_{\pi(\mathbf{x})}\left[\log\frac{\pi(x_{i_{\ell}}|\mathbf{x}_{-i_{\ell}})}{\pi(x_{i_{\ell}})}\right],$
@@
$\mathbb{E}_{\pi(\mathbf{x}_{z_{<k}})}[\mathrm{TC}_{\pi}(z_{k}|\mathbf{x}_{\mathbf{z}_{<k}})]\leq\left(1-\frac{1}{s_{k}}\right)\sum_{i\in z_{k}}\mathbb{E}_{\pi(\mathbf{x})}\left[\log\frac{\pi(x_{i}|\mathbf{x}_{-i})}{\pi(x_{i})}\right].$
From the above, and using $\sum_{k}\sum_{i\in\mathbf{z}_{k}}\mathbb{E}_{\pi(\mathbf{x})}\left[\log\frac{\pi(x_{i}|\mathbf{x}_{-i})}{\pi(x_{i})}\right]=ND(\pi)$ for any partition $\mathbf{z}$, as well as $s_{k}\leq s_{\max}$, we obtain (5). ∎
@@
A direct computation yields $\mathrm{TC}_{\pi}(z_{k}|\mathbf{x}_{\mathbf{z}_{<k}})=(s_{k}-1)\log|\mathcal{X}|$ and $D(\pi)=\log|\mathcal{X}|$. Summing over $k$ gives $E_{\text{fact}}=(N-K)\log|\mathcal{X}|$, i.e. equality in (4) and equality in (5) if $N/s_{\max}=K$. ∎
If $s_{k}$ is roughly constant over $k$, then $s_{\max}\approx N/K$, so that the upper bound in (5) is approximately $(N-K)D(\pi)$. In particular if $K,N\to+\infty$ and $N/K\to\bar{s}>1$, the bound in (4) grows linearly with $N$, resulting in very weak guarantees on the overall sampling error. However, such worst-case bounds can be very pessimistic: below we show that if $\mathbf{z}$ is appropriately randomized, then $E_{\text{fact}}$ is provably of much smaller size.
@@ 4 The random-order case
Consider now the situation where $z_{k}$ is sampled by first generating its size $s_{k}=|z_{k}|$ and then sampling its entries uniformly without replacement from $\{1,\dots,N\}\backslash\mathbf{z}_{<k}$. This is the case in common implementations of MDMs [16]. Specifically, to generate the ordered partition $\mathbf{z}$, the algorithm first sample the sizes $(s_{1},s_{2},\ldots,s_{K})=\mathbf{s}$ with $\sum_{k=1}^{K}s_{k}=N$. Then, once these sizes are sampled, recursively
@@ 4.1 Rewriting the factorization error with the information profile
On the other hand, by (7), if $\mathbf{z}\sim\nu^{\theta}$ and $i\in z_{k}$ uniformly at random then $(i,\mathbf{z}_{<k})$ has distribution equal to $(\sigma_{a_{k-1}+1},\{\sigma_{1},\ldots\sigma_{a_{k-1}}\})$ with $\sigma\sim\mathrm{Unif}$. Thus, since $|z_{k}|=s_{k}=a_{k}-a_{k-1}$, we obtain
@@ 4.2 Resulting upper bounds
A bound similar to the one in Theorem 7 was recently derived in Li and Cai [10, Theorem 1] with a different and significantly less direct proof approach.
@@ 4.3 Explicit computations with geometric schedules and lower bounds
Interestingly, one can compute almost exactly the value of $E_{\text{fact}}$ for the case of random, geometrically distributed schedule sizes, without requiring any assumption on $\pi$. Specifically, given $p\in(0,1)$ and $m\in\mathbb{N}$, let $\mathrm{Geom}(p;m)$ denote a Geometric distribution starting from $1$ and with a threshold at $m$, i.e. a random variable $X\sim\mathrm{Geom}(p;m)$ satisfies $\mathrm{Pr}(X=i)=(1-p)^{i-1}p$ for $i\in\{1,\dots,m-1\}$ and $\mathrm{Pr}(X=m)=(1-p)^{m-1}$. The proof of the following result relies on the memoryless property of the geometric distribution, it can be found in the appendix.
@@
$E_{\text{fact}}\leq\frac{1-p}{p}D(\pi),$
@@
$E_{\text{fact}}\geq\frac{1-p}{p}D(\pi)-\left(\frac{1-p}{p}\right)^{2}\left(\max_{i=1,\dots,N-1}\Delta f(i)\right)\,.$